OpenStudy (surryyy):

Determine the domain of the function f(x) = 2x - 4 when the range is {0, 12, 20}. {2, 8, 12} {0, 12, 20} {4, 16, 24} none of these

Nnesha (nnesha):

domain= x values range = y values

Nnesha (nnesha):

so substitute y for given set {0,12,20)one by one to find domain

Nnesha (nnesha):

how ?

OpenStudy (anonymous):

you need to solve for them and not guess

Nnesha (nnesha):

f(x) is same as y so \[\huge\rm y=2x-4\] start with first number 0 substitute x for 0 what do you get ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha f(x) is same as y so \[\huge\rm y=2x-4\] start with first number 0 substitute x for 0 what do you get ? \(\color{blue}{\text{End of Quote}}\) i meant substitute y**8

OpenStudy (anonymous):

no no no you do not substitute x for 0 you substitute f(x) to each the values in the range. you said it yourself y = f(x) range = y domain = x so, \(f(x) = 2x - 4 \) means solve for x for each of the following: \(0 = 2x-4 \) \(12 = 2x-4 \) \(20 = 2x-4 \)

OpenStudy (anonymous):

if you do not know how to solve for x, then you have no business taking the course.

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \(\color{blue}{\text{Originally Posted by}}\) @Nnesha f(x) is same as y so \[\huge\rm y=2x-4\] start with first number 0 substitute x for 0 what do you get ? \(\color{blue}{\text{End of Quote}}\) i meant substitute y** \(\color{blue}{\text{End of Quote}}\) ^^^^^

OpenStudy (surryyy):

x = 2 x=8 x=12

OpenStudy (anonymous):

how did you GET those values? CAN YOU SHOW YOUR SOLUTION FOR EACH?

OpenStudy (surryyy):

0=2x−4 2x−4+4=0+4 2x=4 2x/2 = 4/2 x=2

OpenStudy (surryyy):

12=2x-4 2x−4+4=12+4 2x=16 2x/2 = 16/2 x=8

OpenStudy (surryyy):

20=2x-4 2x−4+4=20+4 2x=24 2x/2 = 24/2 x=12

OpenStudy (surryyy):

happy? fgt.

Nnesha (nnesha):

PERFECT! great job!! :=)

Nnesha (nnesha):

fgt????

OpenStudy (surryyy):

ty Nnesha.

Nnesha (nnesha):

yw